The bottler of a popular soft drink can purchase glass bottles for $0.04 each or all-aluminum cans for $0.05 each. If the aluminum cans are used, she believes she can achieve a saving of $0.005 per can in shipping and handling costs, owing to the reduced weight. In addition, setting up a reclamation drive, she believes she can recover about 50% of the cans by offering to pay $0.10 per pound for returned cans at four return centers which she will establish in the area her company serves. The used cans weigh 20 to the pound, and she can sell the used cans to the manufacturer for $0.15 per pound. She estimates that he overhead cost of operating the return centers will not exceed 20% of the cost of repurchasing the used cans. In addition, she believes that she will obtain favorable response from ecology-minded customers. Which type of beverage container would you recommend that she use?


1
Expert's answer
2021-02-24T10:39:00-0500

Let's consider two options:

Option 1 is to buy glass bottles for $0.04 each.

Option 2 is to buy all-aluminum cans for $0.05 each.

Let the shipping and handling costs for Option 1 be denoted by SHC.

Hence, SHC for option 2 are equal to SHC - $0.005 per unit.


Let's evaluate the profit from reclamation drive.

Let's assume that the N aluminum cans re sold.

Hence, 0.5N cans (50% of the cans sold) can be recovered.

Weight of 20 cans is equal to 1 pound. Hence, 1 can's weight is equal to 0.05 pound.

If she can receive net $0.05 per pound ($0.15 - $0.10), the total revenue from reclamation drive (denoted by TRRD) can be calculated as follows:

"TRRD=0.5\times N\times0.05\times 0.05=0.00125\times N"

Hence, TRRD per unit is equal to $ 0.00125.

The overhead cost of operating the return centers is equal to 20% of the cost of repurchasing the used cans (equal to $0.00125 per unit). Hence, net TRRF per unit can be calculated as follows:

"TRRDnet=0.00125\times N-0.2\times 0.00125\times N=0.001\times N" .

Hence, TRRD per unit is equal to $ 0.001.

Unit total expenditures for one beverage container for Option 1:

SHC+$0.04

Unit total expenditures for one beverage container for Option 2:

SHC-$0.005+$0.05 -$0.001=SHC+$0.044

Conclusion: Option 2 is more expensive for company.

ANSWER: the glass bottles is recommended to use